OpenStudy (anonymous):

graph the equation. identify the conic sequence, any lines of symmetry, and the domain and range

OpenStudy (anonymous):

\[x=2y^2+5\]

OpenStudy (anonymous):

this is similar to the one we just did but alittle confused...... 0.o help @whpalmer4

OpenStudy (anonymous):

ok can the domain and range be \[-5\le x \le 5 \] & \[-5 \le y \le 5\] ????

OpenStudy (whpalmer4):

okay, have you figured out what the graph looks like?

OpenStudy (whpalmer4):

that's going to tell you what conic section, too

OpenStudy (anonymous):

um no i dont know how to graph these??? sorry it took so long for me to get back to you and my computer kept glitching

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Okay, make yourself a table with two columns, x and y. Pick some values for x. Write them in the table. Calculate the values of y that go with them. Write them in the table as well. Now break out the graph paper and plot those points. What does the graph look like?

OpenStudy (anonymous):

ok what plots do i use. please give me a example

OpenStudy (anonymous):

lol points

OpenStudy (whpalmer4):

Our equation is \[x = 2y^2+5\] For this it's easier to use y as the dependent variable, so I'll do that. y = -2, x = 2(-2)^2+5 = y = -1, x = 2(-1)^2+5 = y = 0, x = 2(0)^2+5 = y = 1, x = 2(1)^2+5 = y = 2, x = 2(2)^2+5 =

OpenStudy (anonymous):